Attending: Bruno, Vlad, Alex, Cristiana, Adam, Tapasi

Computing issues

CI login certificates transitioning. For keep-ups this was a hiatus but is now fixed. The dCache had unintended consequences that we still don’t know, so Alex will try and figure out which were the problems and the timeline. Stascache fixing is ongoing. Tape jobs were being held (probably run time limits). We didn’t design the system for this, so it’s not surprising. LEMserver seems to not be working right now. Bruno will try to run tomorrow to figure out the issues.


Dan still can’t log in, despite the fact that he seems to have the right certificate. We need the snapshot, the ana branch and the root 6 branch, so this is top priority. Bruno suggests asking Jonathan.

Scrum board

ND top-up

Local test run successfully

FD CRY with new fibre model

There is a problem with the metadata. The output seems to be wrong, so we need a fix in NGU.

New sample: ND with cosmic overlay

System improvements

Declaring file after copying (Vlad)

He just had to change a few lines. He checked and committed. He also spotted some errors that we didn’t know of and were probably there for a long time. The standard version, and it has the same error. This may be related from a scheme change in SAM, so we may now need to pull the size of the file. We would need to add that to Vladimir will make this fix.

Metadata defaults (Cristiana)

Contact Joseph.

Job-specific metadata (Adam)

He talked to Alex, but there were power problems in Duluth

Including horn current metadata

We need to identify a champion for this project


Modified the certification scheme. A few bumps but it’s now working ok. Changed subrun end runtime to start time.


Enrique has 2 tickets open about problems in Cincinnati and SMU. He will add Alex and Bruno as watchers.